OpenStudy (ulahlynn):

Can someone please help me with Permutation? :) 3 boys and girls are to be photographed. In how many ways can they be arranged in a row: (a) if the girls are photographed by themselves. (b) if the boys are photographed by themselves.

OpenStudy (dragonborn):

ok, so let the boys be ABC and the girls be DEF if u are just looking at the girls, you can do DEF EFD DFE FED FDE EDF, so 6 different ways, and since there are the same amount of boys, it would be the same just different letters

OpenStudy (ulahlynn):

ohhh.. so the answer is 12??

OpenStudy (dragonborn):

o, theres only 2 girls...

OpenStudy (ulahlynn):

and im assigned to answer no. 8 can you help me with the rest. if its ok with you?

OpenStudy (dragonborn):

O.O i would have to work on that at home, im in school right now >_< someone can probrably answer it for u before then

OpenStudy (anonymous):

which one you need help with?

OpenStudy (ulahlynn):

well its ok.. thanks Dragonborn :)

OpenStudy (ulahlynn):

i need help answering a - g can you help me? please

OpenStudy (anonymous):

sure lets do the easy one first two girls, a and b ab, ba two ways to do it

OpenStudy (anonymous):

3 boys, x, y, z xyz xzy yxz yzx zxy zyx six ways to do it

OpenStudy (anonymous):

now you don't want to continue to have to write all the combinations, so we think 3 boys 3 choices for first one 2 for the second 1 for the third by the counting principle the number of ways to do this is \[3\times 2\times 1\] often written as \(3!\)

OpenStudy (ulahlynn):

ohhhh.. haha i think i kinda get it.

OpenStudy (anonymous):

that makes the next one easy 5 people arranged in a row, number of ways to do it is \[5!=5\times 4\times 3\times 2\times 1\]

OpenStudy (anonymous):

now we get to a harder one boy, girl, boy, girl, boy but it is not that hard 3 choices for the first boy 2 choices for the first girl 2 choices for the second boy 1 choice for the second girl 1 choice for the last boh by the counting principle, the number of ways to do this is \[3\times 2\times 2\times 1\times 1\]

OpenStudy (anonymous):

notice that we did not use any formula at all, it was all based on the counting principle, and reasoning how many choices at each step all the formulas are based on that as well

OpenStudy (ulahlynn):

ohhhh....

OpenStudy (ulahlynn):

soo letter g is the answer 24?? im not that sure though

OpenStudy (anonymous):

the row begins and ends with a girl?

OpenStudy (anonymous):



OpenStudy (anonymous):

how many choices for the first girl ?

OpenStudy (ulahlynn):

3 choices for the first boy 2 choices for the first girl 2 choices for the last boy ohh imeant 12 and letter e ! .. hahahhaa sorry

OpenStudy (ulahlynn):

i dunno if its right

OpenStudy (anonymous):

ok now i have to go back and look why don't you do g whilst i check your e ?

OpenStudy (ulahlynn):

ok :)

OpenStudy (anonymous):

you are close withe, but it not quite right, we can fix it though let me know what you get for f, take a look at the diagram i sent

OpenStudy (ulahlynn):

2 x 3 x 2 x 1 x 1 = 12 also ??

OpenStudy (anonymous):

yes, that is right

OpenStudy (anonymous):

that is for f,

OpenStudy (anonymous):

Hi!...I don't mean to interrupt but if you watch this video, I think you might just master permutations! :o) https://www.youtube.com/watch?v=XqQTXW7XfYA

OpenStudy (anonymous):

good work! now we have to redo e more carefully, because it is not 12

OpenStudy (ulahlynn):

(f) if no boy sits next to a girl.

OpenStudy (ulahlynn):

ill try watching it :)) thankyouu @SinginDaCalc2Blues

OpenStudy (anonymous):

oh i got confused, the answer of 12 is for question g not f want to do f next ?

OpenStudy (ulahlynn):

yeah. hahaha

OpenStudy (ulahlynn):

but honestly i dont know how to do F

OpenStudy (anonymous):

i'm all messed up with the letters, so lets just say what the question is

OpenStudy (ulahlynn):

ok.. hahhaha :D

OpenStudy (anonymous):

how can they be arranged if no boy sits next to a girl?

OpenStudy (anonymous):

i would say that was impossible in one row

OpenStudy (ulahlynn):

I Know Right!! like honestly i dont get it

OpenStudy (anonymous):

what is wrong with saying 0 ways?

OpenStudy (ulahlynn):

hahhahaha.. i think its valid.. sooo 0 ways then

OpenStudy (anonymous):

some boy has to sit next to some girl for sure, so zero is a perfectly good answer

OpenStudy (anonymous):

which ones are left?

OpenStudy (ulahlynn):

uhhmm.. is my e alright?? the question is "if the row begins and ends with a boy?"

OpenStudy (ulahlynn):

3 x 1 x 2 x 1 x 2 = 12 ways ???

OpenStudy (anonymous):

B ? ? ? B

OpenStudy (anonymous):

3 for the first by 2 for the last so you had that part correct

OpenStudy (anonymous):

but the middle should be \(3\times 2\times 1\) because you have 3 people to choose from

OpenStudy (anonymous):

\[3\times 3\times 2\times 1\times 2\]

OpenStudy (ulahlynn):

im incorrect.. how come 3 x 2 x 1??

OpenStudy (ulahlynn):

sorry i kinda dont get it. hahahahahaha

OpenStudy (anonymous):

you have 3 choices for the first by right? and 2 choices for the last boy you said that and got that part right

OpenStudy (anonymous):

now that leaves 3 people, (2 girls and a boy) but it doesn't matter what gender they are, they can sit in the middle 3 spots in any order at all

OpenStudy (anonymous):

and the number of ways you can seat three people in three chairs is \(3!\)

OpenStudy (ulahlynn):

ooohh :D soo the answer is 36 ways

OpenStudy (anonymous):

yes, i think so now what is left ?

OpenStudy (ulahlynn):

i think were all done!!!

OpenStudy (ulahlynn):

THANKYOOUUU SOO MUCH @satellite73

OpenStudy (anonymous):

yw hope you learned how to do more if you need to